If the following parabola has a minimum at y=-4 and roots at x=5 and 7. What is the a in the equation ? The equation is y= a(x-p)2+q


Answer by scott8148(6628)

the axis of symmetry is midway between the roots at x=6

the vertex (minimum) is on the axis of symmetry at (6,-4)

y = a(x - 6)^2 - 4 ___ 0 = a(5 - 6)^2 - 4